Homework 6, March 28

Math 402 – Spring Semester, 1990 Second Course in Abstract Algebra, II Prof. Daniel Grayson

2:00 PM MWF, Section F1, 243 Altgeld Hall

Announcement: I will be out of town on April 6, 16, 18, 20, and 27, in order to attend a K-theory converence, an algebraic geometry conference, and a thesis defense. Steve Ullom will take over on April 6 and 27, and Gerald Janusz will take over on April 16, 18, and 20.

Read Section IV.3. These problems are due Wednesday, April 4. Do p. 317: 2, (3), 6, 8. Do p. 325: 1. The exercises in parentheses are to be turned in only if you are uncertain whether you

got the correct answer or whether you could answer it in an oral exam.

The following exam question is to be done independently.

Exam Problem 3: Let F be a field, and let R be the polynomial ring F [x, y] in the indeterminates x and y. Show that not every ideal of R is a free R-module.